OpenStudy (anonymous):

a train leaves a station and travels north at 50 km/hr. two hours later, a second train leaves on a parallel track and travels north at 90 km/hr. how far from the station will they meet ? help PLEASE:D

OpenStudy (amistre64):

if they are on 2 seperate tracks; they never meet

OpenStudy (amistre64):

the idea here is to see when the time involved produces equal distances

OpenStudy (amistre64):

A goes 50 for 2 hours; he got 100 head start. 100 + 50t = some Distance 90t = the same amount of Distance when does: 90t = 100+50t ?

OpenStudy (anonymous):

so i combine like terms there then solve for t? :)

OpenStudy (amistre64):

yep; then use that "t" in either equation to get the distance traveled

OpenStudy (anonymous):

ok i got 4 so now i can multipy by 90t?

OpenStudy (amistre64):

you can, but let me chk the math. 90t = 100 + 50t 9t = 10 + 5t -5t -5t ------------ 4t = 10 t = 10/4 = 5/2

OpenStudy (amistre64):

you got 4, but forgot to divide it off to find "t" all by itself

OpenStudy (amistre64):

or something like that

OpenStudy (amistre64):

90(5)/2 = 45(5) = 200 + 25 = 225

OpenStudy (anonymous):

wow, thanks so much :D

OpenStudy (amistre64):

youre welcome